OpenStudy (anonymous):

For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

@genius12 @ganeshie8 @.Sam. @hartnn @amistre64 someonehelp me please

OpenStudy (anonymous):

Write something like: Michelle is selling lemonade for $2 each and chocolate chip cookies for $1 each. Bob is selling lemonade for $3 each and chocolate chip cookies for $4 each. Michelle makes $9 at the end of the day and Bob makes $26 at the end of the day. If they both sold the same amount of Lemonade and Cookies, how many cookies/lemonade did each person sell? (Note: They can sell cookies/lemonades in parts as well, e.g. michelle can choose to sell 1/2 cup of lemonade for $1 instead of one full cup for $2.)

OpenStudy (anonymous):

Now just solve the system of equations; preferably with 'Elimination' for part b.

OpenStudy (anonymous):

@R00nnYBraiNsbiG

OpenStudy (anonymous):

ok thaks

OpenStudy (anonymous):

ooh wait can u write the equeation please

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

nvm thanks

OpenStudy (anonymous):

@R00nnYBraiNsbiG medal/fan? lol =[

OpenStudy (anonymous):

@genius12 can u help with the elemination

OpenStudy (anonymous):

Ok. So the way elimination works is this: We choose a variable to 'eliminate'. Here we have 2 possible options; x and y. Which one do you want to eliminate?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Ok. So to eliminate y, we must make the coefficient of y (it doesn't matter if the coefficient is negative or positive, as long as it's the same number is all we want at first) the same for both equations. We notice that the y in the second equation has a coefficient of 4 whereas the first one has a coefficient of 1. So to make the coefficients the same, we multiply the entire first equation by 4 to make the y coefficient the same as the second equation:\[\bf \times4 \rightarrow 2x+y=9 \rightarrow 4(2x+y)=4(9) \rightarrow 8x+4y=36\]Now our sytem of equations looks like:\[\bf 8x+4y=36\]\[\bf 3x+4y=26\]To eliminate y, we will subtract the two equations from each other. When we do that, the "4y" will cancel out or get "Eliminated" and we will be left with a number on the right side and with some 'x' on the left side:\[\bf \ \ \ 8x+4y=36\]\[\bf \ \ \ 3x+4y=26\]\[\bf - \_\_\_\_\_\_\_\_\_\_\_\_\_\]\[\bf \ \ \ \ \ 5x=10\]\[\bf \ \ \ \ \ x=2\]Now plug this value of x = 2 back in to any one of the equations and solve for 'y'. And you're done. @R00nnYBraiNsbiG

OpenStudy (anonymous):

OMG ur a life saver

OpenStudy (anonymous):

ok can i get help with one more? @genius12 You graph a system of equations and get two lines that are parallel. Explain what this tells you about the: solution to the system the two equations the result if the system was solved algebraically

OpenStudy (anonymous):

@.Sam. help me please

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

The slope of both lines is the same; there is no solution to the system, hence the system is inconsistent.

OpenStudy (anonymous):

Now I have to finish my AP European History culminating and e-mail it asap to my teacher. It was due almost 2 weeks ago lol and it's the only thing standing in between summer and me so I must work on it. Please don't tag me for the coming couple hours, perhaps after that. @R00nnYBraiNsbiG